The input current of a transformer is 4 A. The primary coil has 50 loops and the secondary has 25 loops. The current that the transformer puts out is
Answer #1

Given
input current of a transformer I1= 4 A.
primary coil N1= 50 loops
secondary coil N2=25 loops
We know transformer ratio
N1/N2=I2/I1
Substituting the value
(50/25)=(I2/4)
I2=8 A

Hope u like it